Tickets for a concert were sold to adults for $3 and to students for $2. If the total receipts were $824 and twice as many adult tickets as student tickets were sold, then how many of each were sold?
Answer by checkley77(12569)

2x*3+x*2=824
6x+2x=824
8x=824
x=824/8
x=103 ans. for the number of student tickets sold.
28103=206 ans. for the number of adult tickets sold.
Proof:
206*3=103*2=824
618+206=824
824=824